Analysts forecast that Michael Hill’s (AU) dividends in 2019, 2020 and 2021 will be the same amount as the dividends paid in 2018 which was 0.05. After 2021 dividends will increase by 2% p.a. What is a fair price for a Michael Hill share in 2019? Required return rate is 9%
Answer #1

Answer:

Value of share in 2021 = Dividend 2022 / (Required return rate - Growth rate) = 0.05 * (1 + 2%) / (9% - 2%) = 0.7285714

Price of share in 2019 = Dividend 2020 / (1 + Required return) + Dividend 2021 / (1 + Required return) 2 + Value of share in 2021 / (1 + Required return) 2

= 0.05 / (1 + 9%) + 0.05 / (1 + 9%) 2 +  0.7285714 / (1 + 9%) 2

= 0.70118

Price of share in 2019 = 0.70118 = 0.70